Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

sibling probability

Expert replies
by dikku07 » Mon Oct 05, 2009 1:38 pm
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
(a) 5/21
(B) 3/7
(C) 4/7
(D) 5/7
(E) 16/21
Join the discussion
Source: — Problem Solving |

Re: sibling probability

by Stuart@KaplanGMAT » Mon Oct 05, 2009 2:31 pm
dikku07 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
(a) 5/21
(B) 3/7
(C) 4/7
(D) 5/7
(E) 16/21
As with most GMAT questions, there are a number of different approaches we can use to solve. Let's look at one of them.

First, let's talk formulas:

Probability = # desired outcomes / total # of possibilities

is the basic probability formula. Chances are pretty good that you use that one already. However, there's another key formula that's a huge time saver on the GMAT:

Prob(want) = 1 - Prob(don't want)

There are a lot of GMAT probability questions on which it's quicker to calculate the probability of what you don't want to happen and subtract that from 1.

On this question, we could certainly take that approach and set up the formula:

Prob(not siblings) = 1 - Prob(are siblings)

Prob(are siblings) = (# of sibling pairs we could possibly pick)/(total # of possible pairs).

Based on the original info, we know that 4 of them have exactly 1 sibling and 3 have exactly 2 siblings. The only way that can be broken down with 7 people is:

1 pair of sibs
1 pair of sibs
1 trio of sibs

So, how many sibling pairs can we make?

From each pair of sibs we can, not surprisingly, make 1 pair of sibs... so that's 2.

From the trio of sibs, we can make 3C2 ("three choose two" is how we pronounce that in the combinatorics world) = 3!/2!1! = 3 pairs of sibs. Of course, you really don't need a fancy formula to figure that out - you can just think "If I have 3 people A, B and C, how many pairs can I make? AB/AC/BC... that's 3 pairs."

So, we have 5 different sibling pairs.

For the total number of possible pairs, we do want to use the combinations formula:

7C2 = 7!/2!5! = 7*6/2 = 21

Therefore, the probability of choosing a pair of siblings is: 5/21

Is that one of the answers? Of course it is. The GMAT knows that a large number of test takers will be so happy that they completed a calculation that they'll forget what the original question was! For just that reason, we have Step 4 of the Kaplan Method for Problem Solving: double check the question.

We're careful, so we realize that the correct answer to the original question isn't 5/21, it's:

1 - 5/21 = 16/21... choose (E).
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by dikku07 » Mon Oct 05, 2009 2:59 pm
Thanks Stuart for in depth explanation. thats correct answer
Join the discussion